OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (anonymous):

well @jose7895 did you solve your last question?

OpenStudy (anonymous):

i helped you how to solve.you can use it here and solve this question.

OpenStudy (anonymous):

can you help me solve this one as well? @PFEH.1999

OpenStudy (anonymous):

well,you have x = -4 so what is (2)x^2 ?

OpenStudy (anonymous):

first find x^2 and then you can find 2(x^2)

OpenStudy (anonymous):

2(x^2)=-16

OpenStudy (anonymous):

but x= -4 so x^2 = (-4)(-4) = 16 and 2x^2 is 32

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sty i have to go you can do it for the rest and find the values: 2x2 2z4 y2 x2 z4 and then you can solve it easily

OpenStudy (anonymous):

2x2 = 32 2z4 = (2)2^4 = 64 y2 = 3^2 = (3)(3) = 9 can u do the rest?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

well you can give me a medal by clicking on Best Response button adn close the question as you the answer ;)